# Can Vehicle Motion Planning Generalize to Realistic Long-tail Scenarios?

### v1

@@ I Introduction
The interPlan benchmark. During complex traffic maneuvers there is often mutual influence between traffic participants. Mastering these bi-directional interactions is crucial for smooth progress, ultimately making it a fundamental enabler for autonomy [3, 4]. Therefore, we propose a novel benchmark to evaluate vehicle motion planning methods in highly interactive and rare scenarios. We take nuPlan [1] scenarios as a starting point for our benchmark and augment them with additional agents, obstacles, or alternative navigation goals, creating realistic long-tail scenarios. In total, our benchmark comprises 80 scenarios covering the following situations: Nudging around parked vehicles, overtaking obstacles, passing a construction zone, passing an accident site, facing jaywalkers, as well as lane changes in low, medium, and high traffic density.
@@
  • We release code for both the interPlan benchmark as well as integration of LLMs with the nuPlan framework, hoping to facilitate further research on the topic.
@@ III-B Scenario Generation
Realistic long-tail scenarios. In our benchmark, we aim to test long-tail scenarios, which are underrepresented in real-world driving and thus might not occur in large-scale training data. Such scenarios include encountering a lane blocked by construction zones (cf. Fig. 0(a)), nudging around (cf. Fig. 0(b)) or overtaking (cf. Fig. 0(c)) a parked vehicle, and facing jaywalkers at bus stops (cf. Fig. 0(d)). We generate these scenarios for our benchmark by adding objects such as traffic cones, parked vehicles, or stopped busses to the original nuPlan scenarios. Moreover, pedestrians who start crossing the street right when the EV approaches test the planner’s ability to anticipate such non-compliant behaviors. In addition, we include scenarios where the planner has to navigate around an accident site (cf. Fig. 0(e)). This is particularly challenging because stationary touching vehicles on intersections or lanes are rarely if at all, found in real-world datasets. Thus, sophisticated reasoning is required to understand that these vehicles have collided and cannot be expected to start moving, so they must be overtaken carefully. We design the crash sites in such a way that they represent common accidents such as rear-end collisions or collisions with crossing traffic.
@@
Interactive lane-change scenarios. A major shortcoming of the original nuPlan benchmark is that it can be solved almost perfectly by simple lane-following, thus with little interaction with surrounding traffic. For instance, the challenge-winning method PDM-Closed [2] is by design unable to perform lane changes but achieves a nearly perfect score. To counteract this, we incentivize lane changes by augmenting the goal of original nuPlan scenarios. For instance, in the scenario shown in Fig. 0(f), the EV has to do 3 lane changes. Moreover, we initialize the scenarios with different levels of traffic densities, i.e., low (LTD), medium (MTD), and high traffic density (HTD), by randomly spawning traffic agents around the EV.
@@
Augmented traffic agent behavior. Real-world drivers exhibit diverse behaviors ranging from conservative and careful driving to assertive and reckless maneuvers. Planning methods have to be able to handle these varying conditions, especially without knowing what they are about to encounter. Our benchmark tests the ability of planning methods to make safe decisions under these varying conditions by applying different policies to control the surrounding traffic agents. Hence, in addition to nuPlan’s default reactive IDM policy, which breaks as soon as the EV enters its lane, we implement an assertive policy, which only reacts to the EV when it has fully merged into the lane. Simulation of merging scenarios can be carried out with all agents being controlled by the conservative reactive policy, the assertive policy, or a mixed policy, which randomly assigns a policy to each agent. The lane-change scenarios in our benchmark combine these traffic agent policies with different traffic densities. Thus, lane changes in HTD require complex interaction and potentially even prioritizing safety over progress. At the same time, LTD scenarios test a planner’s general ability to merge into another lane.
@@ III-E Comparison to Val14 Mining
Method Val14 $\uparrow$ interPlan normal-$\uparrow$ Constr. $\uparrow$ Acc. $\uparrow$ Jayw. $\uparrow$ Nudge $\uparrow$ Overt. $\uparrow$ LTD $\uparrow$ MTD $\uparrow$ HTD $\uparrow$ Driv. $\uparrow$ Goal $\uparrow$ No-Col. $\uparrow$
learned
Urban Driver [28] 50 4 0 0 0 0 0 0 29 0 30 30 47
@@ IV-A Generalization of state-of-the-art Planning Methods
Our experimental results are shown in Tab. III. We find that PDM-Closed achieves the best results with an overall score of 42%. However, in contrast to previous results on the Val14 test-split and the nuPlan leaderboard, it is far from achieving a perfect score. In particular, it achieves the best score among all state-of-the-art planners in the Construction Zone, Overtake, and Nudge scenarios. This is mainly due to its ability to sample different lateral offsets from the centerline it is following, enabling it to pass some of the obstacles in the lane. Moreover, it outperforms all other methods in the LTD, MTD, and HTD lane change scenarios. However, due to its limitation of not being able to do lane changes, this merely reflects its ability to find an optimal trajectory concerning comfort and safety since it is unable to reach the goal lane in these scenarios. In addition, almost all methods in our benchmark fail in the Acc. and Overt. scenarios, resulting in a score of 0%. Exceptions are DTPP for Acc. and PDM-Open and PDM-Closed for the Overt. scenarios. While DTPP successfully makes a lane change before the stationary vehicles in the Acc. scenario shown in Fig. 1(a), PDM-Closed stops before them (see  1(e)). Similarly, PDM-Open and PDM-Closed pass the parked car in the easiest of the 10 overtake scenarios, i.e., the one with no oncoming traffic. For PDM-Open, this is not intentional but instead caused by a failure to follow the curved centerline since this model does not consider surrounding vehicles. Surprisingly, PDM-Closed performs worse than IDM in the scenarios with Jaywalkers even though it employs an IDM policy for centerline following. In contrast to IDM, PDM uses a future horizon of only 2.0s to evaluate potential longitudinal profiles. In two scenarios, a collision is unavoidable when the pedestrian is inside this horizon. In a further scenario, PDM-Closed avoids a collision by evading the pedestrians at high speed (see Fig. 1(f) instead of stopping and waiting for them to cross (see Fig. 1(b)). Overall, our results reveal critical shortcomings in the generalization capabilities of state-of-the-art planning methods.
@@ V-A LLM Waypoints Planning
GPTDriver [6] fine-tunes and instructs an LLM to predict a set of future waypoints. The prompt comprises general task instruction, perception context, and ego-states. Chain-of-thought reasoning is used to improve the interpretability and quality of the output. We adapt the method to the nuPlan requirements by changing the output to an 8-second trajectory at 2 Hz and by including route information into the prompt. Thus, we provide the planner with a ‘Mission Goal‘ telling it which lanes are on route. We test various LLMs, including Llama-7B, Llama-13B [56], and GPT-3.5. We call this baseline LLMWaypointsPlanner and fine-tune it on a set of 600 diverse and unaugmented nuPlan scenarios for 3 epochs. For Llama we employ QLoRa [57] using a rank of 64 and a learning rate of 2e-4. For GPT, we use openAI’s fine-tuning API.
@@ V-B LLM Behavior Planning
Model LLM interPlan normal-$\uparrow$ Constr. $\uparrow$ Acc. $\uparrow$ Jayw. $\uparrow$ Nudge $\uparrow$ Overt. $\uparrow$ LTD $\uparrow$ MTD $\uparrow$ HTD $\uparrow$ Driv. $\uparrow$ Goal $\uparrow$ No-Col. $\uparrow$

### v2

@@ I Introduction
The interPlan benchmark. During complex traffic maneuvers there is often mutual influence between traffic participants. Mastering these bi-directional interactions is crucial for smooth progress, ultimately making it a fundamental enabler for autonomy [3, 4]. Therefore, we propose a novel benchmark to evaluate vehicle motion planning methods in highly interactive and rare scenarios. We take nuPlan [1] scenarios as a starting point for our benchmark and augment them with additional agents, obstacles, or alternative navigation goals, creating realistic long-tail scenarios. In total, our benchmark comprises 80 scenarios covering the following situations: Nudging around parked vehicles, overtaking obstacles, passing a construction zone, passing an accident site, facing jaywalkers, as well as lane changes in low, medium, and high traffic density. A comparative overview is shown in Tab. I.
@@ III-B Scenario Generation
Realistic long-tail scenarios. In our benchmark, we aim to test long-tail scenarios, which are underrepresented in real-world driving and thus might not occur in large-scale training data. Such scenarios include encountering a lane blocked by construction zones (cf. Fig. 1(a)), nudging around (cf. Fig. 1(b)) or overtaking (cf. Fig. 1(c)) a parked vehicle, and facing jaywalkers at bus stops (cf. Fig. 1(d)). We generate these scenarios for our benchmark by adding objects such as traffic cones, parked vehicles, or stopped busses to the original nuPlan scenarios while leaving the EV state unchanged. Moreover, pedestrians who start crossing the street right when the EV approaches test the planner’s ability to anticipate such non-compliant behaviors. In addition, we include scenarios where the planner has to navigate around an accident site (cf. Fig. 1(e)). This is particularly challenging because crashed vehicles with intersecting bounding boxes are found rarely, if at all, in real-world datasets. Thus, sophisticated reasoning is required to understand that these vehicles have collided and cannot be expected to start moving, so they must be overtaken carefully. We design the crash sites in such a way that they represent common accidents such as rear-end collisions or collisions with crossing traffic.
@@
Interactive lane-change scenarios. A major shortcoming of the original nuPlan benchmark is that it can be solved almost perfectly by simple lane-following, thus with little interaction with surrounding traffic. For instance, the challenge-winning method PDM-Closed [2] is by design unable to perform lane changes but achieves a nearly perfect score. To counteract this, we incentivize lane changes by augmenting the goal of original nuPlan scenarios. For instance, in the scenario shown in Fig. 1(f), the EV has to do 3 lane changes. Moreover, we initialize the scenarios with different levels of traffic densities, i.e., low (LTD), medium (MTD), and high traffic density (HTD), by spawning traffic agents around the EV. Agents are randomly placed along each lane with sufficient spacing to avoid collisions, with maximum distances of 33m, 50m, and 100m for HTD, MTD, and LTD.
@@
Augmented traffic agent behavior. Real-world drivers exhibit diverse behaviors ranging from conservative and careful driving to assertive and reckless maneuvers. Planning methods have to be able to handle these varying conditions, especially without knowing what they are about to encounter. Our benchmark tests the ability of planning methods to make safe decisions under these varying conditions by applying different policies to control the surrounding traffic agents. Our primary objective is not realistic traffic simulation, which is an ongoing research field itself, but to create diverse and reasonable traffic agents that can effectively test a planner’s interactive behavior. Hence, in addition to nuPlan’s default reactive IDM policy, which breaks as soon as the EV enters its lane, we implement an assertive policy, which only reacts to the EV when it has fully merged into the lane. Simulation of merging scenarios can be carried out with all agents being controlled by the conservative reactive policy, the assertive policy, or a mixed policy, which randomly assigns a policy to each agent. The lane-change scenarios in our benchmark combine these traffic agent policies with different traffic densities. Thus, lane changes in HTD require complex interaction and potentially even prioritizing safety over progress. At the same time, LTD scenarios test a planner’s general ability to merge into another lane.
@@ III-E Comparison to Val14 Mining
Method Val14 $\uparrow$ interPlan $\uparrow$ Constr. $\uparrow$ Acc. $\uparrow$ Jayw. $\uparrow$ Nudge $\uparrow$ Overt. $\uparrow$ LTD $\uparrow$ MTD $\uparrow$ HTD $\uparrow$ Driv. $\uparrow$ Goal $\uparrow$ No-Col. $\uparrow$
learned Urban Driver [28] 50 4 0 0 0 0 0 0 29 0 30 30 47
@@ IV-A Generalization of state-of-the-art Planning Methods
Our experimental results are shown in Tab. III. We find that PDM-Closed achieves the best results with an overall score of 42%. However, in contrast to previous results on the Val14 test-split and the nuPlan leaderboard, it is far from achieving a perfect score. In particular, it achieves the best score among all state-of-the-art planners in the Construction Zone, Overtake, and Nudge scenarios. This is mainly due to its ability to sample different lateral offsets from the centerline it is following, enabling it to pass some of the obstacles in the lane. Moreover, it outperforms all other methods in the LTD, MTD, and HTD lane change scenarios. However, due to its limitation of not being able to do lane changes, this merely reflects its ability to find an optimal trajectory concerning comfort and safety since it is unable to reach the goal lane in these scenarios. In addition, almost all methods in our benchmark fail in the Acc. and Overt. scenarios, resulting in a score of 0%. Exceptions are DTPP for Acc. and PDM-Open and PDM-Closed for the Overt. scenarios. While DTPP successfully makes a lane change before the stationary vehicles in the Acc. scenario shown in Fig. 2(a), PDM-Closed stops before them (see  2(e)). Similarly, PDM-Open and PDM-Closed pass the parked car in the easiest of the 10 overtake scenarios, i.e., the one with no oncoming traffic. For PDM-Open, this is not intentional but instead caused by a failure to follow the curved centerline since this model does not consider surrounding vehicles. Surprisingly, PDM-Closed performs worse than IDM in the scenarios with Jaywalkers even though it employs an IDM policy for centerline following. In contrast to IDM, PDM uses a future horizon of only 2.0s to evaluate potential longitudinal profiles. In two scenarios, a collision is unavoidable when the pedestrian is inside this horizon. In a further scenario, PDM-Closed avoids a collision by evading the pedestrians at high speed (see Fig. 2(f) instead of stopping and waiting for them to cross (see Fig. 2(b)). Overall, our results reveal critical shortcomings in the generalization capabilities of state-of-the-art planning methods.
@@ V-A LLM Waypoints Planning
GPTDriver [6] fine-tunes and instructs an LLM to predict a set of future waypoints. The prompt comprises general task instruction, perception context, and ego-states. Chain-of-thought reasoning is used to improve the interpretability and quality of the output. We apply the method with the same prompt after applying minor changes to adapt it to the nuPlan requirements, i.e., we change the output to an 8-second trajectory at 2 Hz and include route information into the prompt. Thus, we provide the planner with a ‘Mission Goal‘ describing which lanes are on route. We test various LLMs, including Llama-7B, Llama-13B [56], and GPT-3.5. We call this baseline LLMWaypointsPlanner and fine-tune it on a set of 600 diverse and unaugmented nuPlan scenarios for 3 epochs. For Llama we employ QLoRa [57] using a rank of 64 and a learning rate of 2e-4. For GPT, we use openAI’s fine-tuning API.
@@ V-B LLM Behavior Planning
Model LLM interPlan $\uparrow$ Constr. $\uparrow$ Acc. $\uparrow$ Jayw. $\uparrow$ Nudge $\uparrow$ Overt. $\uparrow$ LTD $\uparrow$ MTD $\uparrow$ HTD $\uparrow$ Driv. $\uparrow$ Goal $\uparrow$ No-Col. $\uparrow$
@@ VI Conclusion
VII Limitations
While interPlan extends the widely adopted nuPlan simulator with challenging scenarios, several aspects remain unaddressed. As all scenarios are hand-designed or generated based on heuristics, the benchmark comprises only a limited number of scenarios. Nonetheless, interPlan is able to reveal critical shortcomings in current state-of-the-art methods. Further, interPlan scenarios don’t provide ground truth and thus cannot be used for training. Therefore, planners can only be trained on the regular nuPlan scenarios before testing their generalization capabilities in interPlan. Finally, we believe that data-driven traffic simulation could further increase traffic diversity and make the benchmark more challenging.